OpenStudy (anonymous):

Write an equation in point-slope form for the line with the given slope that contains the given point. 5. slope =-1/2 ; (5,-3)

OpenStudy (anonymous):

@jim_thompson5910

pooja195 (pooja195):

so y1 - y2=m(Xand X2) right? so plug your numbers in

pooja195 (pooja195):

do u know how to do that?

OpenStudy (anonymous):

yes i think so

pooja195 (pooja195):

ok y don't u write the equation then type it

OpenStudy (anonymous):

is it y= -1/2 ?

OpenStudy (anonymous):

\[\huge y-\color{red}{y_1}=\color{blue}m(x-\color{green}{x_1})\]

OpenStudy (anonymous):

\[\color{blue}m=\color{blue}{-\frac{1}{2}}\]

pooja195 (pooja195):

Y1=-3

OpenStudy (anonymous):

\[\color{green}{x_1}=\color{green}5\]

OpenStudy (anonymous):

\[\color{red}{y_1}=\color{red}{-3}\]

OpenStudy (anonymous):

okay so y--3 =-1/2 (x-5?

pooja195 (pooja195):

yes

OpenStudy (anonymous):

thank you :)